OpenStudy (anonymous):

Determine whether the polynomial below can be factored into perfect squares. If so, factor the polynomial. Otherwise, select that it cannot be factored into a perfect square. 16x^2 - 64x+64

OpenStudy (zale101):

16x^2 - 64x+64 Factor out what's common.

OpenStudy (anonymous):

64? so. 16x^2 -x?

OpenStudy (zale101):

16 is the common factor, right?

OpenStudy (zale101):

16x^2 - 64x+64=0 16(x^2 - 4x+4)=0

OpenStudy (zale101):

(x^2 - 4x+4) can you factor this further?

OpenStudy (zale101):

Once you factor, you can know whether it can factor into perfect squares or not.

OpenStudy (anonymous):

It can't

OpenStudy (zale101):

(a+b)^2= a^2+2ab+b^2 (a-b)^2=a^2-2ab+b^2 both of these are perfect squares, you have, can this be factored into a perfect square? (x^2 - 4x+4) You can factor this, try it.

OpenStudy (zale101):

Btw, is this your first time learning about factoring?